What is the average speed of traveling 700 miles in 9 hours?
Found 2 solutions by rajagopalan, rwm:
Answer by rajagopalan(174)

Distance= 700 miles
Time= 9 hours
Speed=D/T
=700/9
=77.77 mph
Average speed = 77.77 miles per hour

Answer by rwm(914)

r*t=d
r*9=700
r=700/9